Find the inverses of the following functions.
f(x)=x^3 +5
and f(x)= 6/7 +7
can someone help me with these I dont understand them can someone explain them to me.

Answer by stanbon(75887)

Find the inverses of the following functions.
f(x)=x^3 +5
1st:: Interchange x and y
x = y^3+5
2nd:: Solve for "y":
y = cbrt(x-5)
-------------------------
f(x)= (6/7)x +7
1st: Interchange x and y
x = (6/7)y + 7
2nd: Solve for "y":
y = (x-7)/(6/7)
y = (7/6)(x-7)
===================
Cheers,
Stan H.
===================